OpenStudy (anonymous):

Im working with Exponential growth and Decay. (logs&lns). A furniture store is closing ouut its buisness. Each week the owner lowers prices by 25%. After how many weeks will the sale price of a %500 item drop below $100?

OpenStudy (anonymous):

100=500(1-(0.25/4)^1/4w 100/500=(1-(0.25/4)^1/4w w=((log100/500)/(log1-(0.25/4))/4 w=6.234 take 6 weeks (log 100/500)/(log

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

nice pic